OpenStudy (anonymous):

find integral of {sec(x)tan(x)/sec(x)-1

OpenStudy (yuki):

could you write it out using the equation button?

OpenStudy (anonymous):

\[\int\limits \frac{\sec(x) \tan(x)}{\sec(x) -1} dx\]?

OpenStudy (anonymous):

so du = sec(x)tan(x) ^_^ and you'll get the following: \[\int\limits \frac{1}{u} du = \ln|u| + c = \ln|\sec(x)-1| + c\]let u = sec(x) -1 Correct me if I'm wrong ^_^

OpenStudy (anonymous):

oh dear, the steps are messed up , I apologize

OpenStudy (anonymous):

thank you im trying to figure it out ;-)

OpenStudy (anonymous):

oh, you're welcome, let me know if you don't get it :)